# Robust Blockwise Random Pivoting:

### v3

@@ Abstract
The interpolative decomposition (ID) aims to construct a low-rank approximation formed by a basis consisting of row/column skeletons in the original matrix and a corresponding interpolation matrix. This work explores fast and accurate ID algorithms from five essential perspectives for empirical performance: (a) skeleton complexity that measures the minimum possible ID rank for a given low-rank approximation error, (b) asymptotic complexity in FLOPs, (c) parallelizability of the computational bottleneck as matrix-matrix multiplications, (d) error-revealing property that enables automatic rank detection for given error tolerances without prior knowledge of target ranks, (e) ID-revealing property that ensures efficient construction of the optimal interpolation matrix after selecting the skeletons. While a broad spectrum of algorithms have been developed to optimize parts of the aforementioned perspectives, practical ID algorithms proficient in all perspectives remain absent. To fill in the gap, we introduce robust blockwise random pivoting (RBRP) that is parallelizable, error-revealing, and exactly ID-revealing, with comparable skeleton and asymptotic complexities to the best existing ID algorithms in practice. Through extensive numerical experiments on various synthetic and natural datasets, we demonstrate the appealing empirical performance of RBRP from the five perspectives above, as well as the robustness of RBRP to adversarial inputs.
@@
Randomized Numerical Linear Algebra, Interpolative Decomposition, Column Subset Selection, Adaptive Sampling, Parallelization
@@ 1 Introduction
The interpolative decomposition (ID) is a special type of low-rank approximation where a basis for the row/column space is formed by subselecting rows/columns from the target matrix. This allows the ID to capture the essential structures and relationships within the data. ID has found applications in various fields, including numerical analysis, data compression, and machine learning, where preserving the integrity of the original data is of paramount importance.
Given a data matrix $\mathbf{X}=\left[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\right]^{T}\in\mathbb{R}^{n\times d}$ consisting of $n$ data points $\left\{\mathbf{x}_{i}\in\mathbb{R}^{d}\right\}_{i\in[n]}$, along with a constant $\epsilon>0$ and a target rank $r\in\mathbb{N}$, we aim to construct an $(r,\epsilon)$-ID of $\mathbf{X}$:
(1) $\underset{n\times d}{\mathbf{X}}\approx\underset{n\times k}{\mathbf{W}}\ \underset{k\times d}{\mathbf{X}_{S}}\quad\emph{s.t.}\quad\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}\leq(1+\epsilon)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}$
where (i) $k$is the (unknown) target rank; (ii) ${\mathbf{X}}_{\langle r\rangle}$is the optimal rank-$r$ approximation of $\mathbf{X}$ (given by the singular value decomposition (SVD)); (iii) $S=\left\{s_{1},\cdots,s_{k}\right\}\subset[n]$are the skeleton indices corresponding to the row skeleton submatrix $\mathbf{X}_{S}=[\mathbf{x}_{s_{1}},\cdots,\mathbf{x}_{s_{k}}]^{\top}\in\mathbb{R}^{k\times d}$; and (iv) $\mathbf{W}\in\mathbb{R}^{n\times k}$is an interpolation matrix with (nearly) optimal error $\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}$ for the given $\mathbf{X}_{S}$. The construction of an ID can be divided into two stages:
@@
    Given $S$, compute an interpolation matrix $\mathbf{W}\in\mathbb{R}^{n\times\left|S\right|}$ efficiently111 It is worth clarifying that although Equation 2 directly provides $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}$ as the minimizer of $\min_{\mathbf{W}}\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$, the explicit computation of $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ takes $O\left(ndk+dk^{2}+nk^{2}\right)$ time (e.g. via Equation 8), where the dominant cost $O(ndk)$ is undesired. Therefore, whenever available, information from the skeleton selection process is usually leveraged to evaluate/approximate $\mathbf{W}$ efficiently in $o(ndk)$ time (as elaborated in Section 5), which may lead to the gap $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)>0$. that (approximately) minimizes the interpolation error: $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)\triangleq\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}$.
In this work, we explore fast and accurate ID algorithms from the perspectives of both skeleton subset selection and interpolation matrix construction, and we start by posing the main question:
How do we construct an $(r,\epsilon)$-ID efficiently with as small a skeleton $S$ as possible,
@@ 1.1 What are Fast and Accurate ID Algorithms?
    Skeleton complexity measures the minimum size of the skeleton set (i.e. the minimum possible rank of ID) that an algorithm must select before $\mathcal{E}_{\mathbf{X}}\left(S\right)\leq(1+\epsilon)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}$ (i.e. where there exists an $(r,\epsilon)$-ID associated with $S$).
@@
    Parallelizability (in the context of this work) refers to the implementation advantage that the dominant cost for skeleton selection in the algorithm appears as matrix-matrix, instead of matrix-vector, multiplications. On modern processors, Level-3 BLAS, or matrix-matrix, operations can leverage memory cashing much more efficiently than Level-2 or Level-1 BLAS operations [3, 23]. As a result, a sequence of matrix operations implemented using Level-3 BLAS can attain significantly higher performance than an optimal implementation using Level-2 or Level-1 BLAS, especially on parallel processors.
@@
    Error-revealing property refers to the ability of an algorithm to evaluate the skeletonization error efficiently on the fly so that the target rank $k$ does not need to be given a priori.
    Definition 1.1 (Error-revealing).
    An ID algorithm is error-revealing if after selecting a skeleton subset $S$, it can evaluate $\mathcal{E}_{\mathbf{X}}\left(S\right)$ efficiently with at most $O\left(n\right)$ operations.
    That is, given any relative error tolerance $\tau\in(0,1)$, an error-revealing ID algorithm can determine an appropriate rank $k$ with negligible additional cost (i.e. no more than $O(nk)$) while selecting a skeleton subset $S$ ($\left|S\right|=k$) such that $\mathcal{E}_{\mathbf{X}}\left(S\right)\leq\tau\left\|\mathbf{X}\right\|_{F}^{2}$. For the sake of analysis, we define a useful constant associated with $\mathbf{X}$ throughout the paper:
    (3) $\eta_{r}\triangleq{\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}}/{\left\|\mathbf{X}\right\|_{F}^{2}},$
    which quantifies the relative tail weight of $\mathbf{X}$ with respect to the rank $r$ and represents the relative optimal rank-$r$ approximation error of $\mathbf{X}$. Notice that when $\tau<(1+\epsilon)\eta_{r}$, the algorithm outputs an $(r,\epsilon)$-ID.
@@
    ID-revealing property characterizes whether the skeleton selection stage of an ID algorithm generates sufficient information for efficient interpolation matrix construction.
    Definition 1.2 (Exactly, approximately, and non-ID-revealing).
    We say a skeleton selection algorithm is exactly ID-revealing if it contains sufficient information in addition to $S$ such that the optimal interpolation matrix $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be evaluated efficiently in $O(nk^{2})$ time. Otherwise, if a suboptimal interpolation matrix $\mathbf{W}\approx\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be constructed in $O(nk^{2})$ time, we say that the skeleton selection algorithm is approximately ID-revealing. If neither $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ nor its approximations can be constructed in $O(nk^{2})$ time, the skeleton selection algorithm is non-ID-revealing.
@@ 1.2 How to Combine Adaptiveness and Randomness for ID?
Adaptiveness and randomness are two critical algorithmic properties that can be leveraged to characterize a skeleton selection algorithm. On the one hand, with adaptiveness, the skeleton selection in each step is aware of the selected skeletons from the previous steps so that redundant skeleton selection can be better avoided. On the other hand, randomness diversifies the skeleton selection, thereby improving the robustness of algorithms to scarce adversarial inputs and providing strong statistical guarantees for skeleton complexities.
@@
    Greedy pivoting is a classical strategy in numerical linear algebra that involves only adaptiveness. As an example of greedy pivoting, column-pivoted QR (CPQR) [22, Section 5.4.2] picks the point in the residual with the maximum norm in each step and adaptively updates the residual via projection onto the orthogonal complement of the skeletons. Despite the error-revealing and exactly ID-revealing abilities, along with the remarkable empirical success, of CPQR, deterministic greedy pivoting algorithms like CPQR are inherently sequential and vulnerable to adversarial inputs where the skeleton complexity approaches the problem size $n$ [29, 5].
@@
    Sampling comprises the well-researched body of skeleton selection methods that involve only randomness. Some common examples related to ID include squared-norm sampling [20], leverage score sampling [33], and DPP sampling [2, 31, 12]. As a trade-off between the skeleton complexity and efficiency, the fast ($O(nd)$-time) sampling methods like squared-norm sampling tend to suffer from high skeleton complexities that depend heavily on the matrix [13, 5]; whereas constructions of the more sophisticated distributions like leverage score and DPP are usually expensive [18, 27, 10]. Moreover, for ID, sampling methods generally fail to be error-revealing or ID-revealing.
@@
    Random pivoting combines adaptiveness and randomness by replacing the greedy selection of maximum norm in CPQR with random sampling according to the squared-norm distribution associated with the residual [13, 14, 5]. Closely related to the ID problem considered in this work, the idea of random pivoting has been revitalized by the inspiring recent work [5] in the context of column Nyström approximation with a nearly optimal skeleton complexity guarantee [5, Theorem 3.1]. However, analogous to CPQR, although random pivoting enjoys the desired error-revealing and exact ID-revealing properties, the sequential nature of random pivoting compromises its efficiency in practice.
@@
    Sketchy pivoting is an alternative combination of adaptiveness and randomness that randomly embeds the row space of $\mathbf{X}$ via sketching [26, 49], on top of which skeletons are selected via greedy pivoting [48, 16, 39]. In contrast to random pivoting, the cost of sketchy pivoting is dominated by the embarrassingly parallelizable sketching process. As a trade-off for superior empirical efficiency, sketchy pivoting sacrifices the error-revealing property since sketching requires prior knowledge of the target $k$ or its overestimation. Furthermore, sketchy pivoting is approximately ID-revealing due to the loss of information during sketching. As a simple effective remedy for the loss of accuracy, we observe that multiplicative oversampling can remarkably improve the gap between the interpolation and skeletonization error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)$, which we refer to as oversampled sketchy ID (OSID).
@@
Table 1: Summary of performance for skeleton selection methods that leverage adaptiveness and/or randomness. Recall from Equation 3 that $\eta_{r}\triangleq{\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}}/{\left\|\mathbf{X}\right\|_{F}^{2}}$. The skeleton complexities in $(\cdot)^{*}$ are conjectured based on extensive empirical evidence and intuitive rationale but without formal proofs. In the “Dominant Cost” column (i.e. showing both the asymptotic complexity and parallelizability), “m-v” stands for matrix-vector multiplications, which are significantly less efficient than “m-m”—matrix-matrix multiplications—due to the lack of parallelizability [23].
@@
Skeleton Selection Skeleton Complexity Dominant Cost Error-revealing ID-revealing Greedy pivoting/CPQR ([22, Section 5.4.2]) $k\geq\left(1-\left(1+\epsilon\right)\eta_{r}\right)n$ $O(ndk)$ m-v Yes Exactly Squared-norm sampling (Proposition 2.3 [20]) $k\geq\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ $O(nd)$ m-m No Non Random pivoting (Algorithm 1 [5, 13, 14]) $k\geq k_{\text{RP}}=\frac{r}{\epsilon}+$ $r\min\left\{\log\left(\frac{1}{\epsilon\eta_{r}}\right),1+\log\left(\frac{2^{r}}{\epsilon}\right)\right\}$ $O(ndk)$ m-v Yes Exactly Sketchy pivoting (Algorithm 3 [48, 16]) $\left(k\gtrsim k_{\text{RP}}\right)^{*}$ $O(ndk)$ m-m No Approximately RBRP (this paper) (Algorithm 4) $\left(k\gtrsim k_{\text{RP}}\right)^{*}$ $O(ndk)$ m-m Yes Exactly
@@
To fill in a critical missing piece in the family of existing methods in Table 1 that leverage adaptiveness and randomness, in this work, we introduce robust blockwise random pivoting (RBRP, Algorithm 4)—an ID algorithm that is $O(ndk)$-time, parallelizable, error-revealing, and exactly ID-revealing, with comparable skeleton complexities to the nearly optimal random pivoting [5] in practice.
@@
In particular, the plain blockwise random pivoting (e.g. an extension of the blocked RPCholesky algorithm introduced in [5, Algorithm 2.2]) tends to suffer from unnecessarily large skeleton complexity (cf. Figure 1 (left)) under adversarial inputs (e.g. Example 4.1) due to the lack of local adaptiveness within each block. As a remedy, RBRP leverages robust blockwise filtering (Remark 4.2)—applying CPQR to every small sampled block locally and discarding the potentially redundant points through a truncation on the relative residual of the CPQR. By choosing a reasonable block size, such robust blockwise filtering effectively resolves the inefficiency in skeleton complexity encountered by the plain blockwise random pivoting (cf. Figure 1 (right)), with negligible additional cost.

### v5

@@ Abstract
The interpolative decomposition (ID) aims to construct a low-rank approximation formed by a basis consisting of row/column skeletons in the original matrix and a corresponding interpolation matrix. This work explores fast and accurate ID algorithms from comprehensive perspectives for empirical performance, including accuracy in both skeleton selection and interpolation matrix construction, efficiency in terms of asymptotic complexity and hardware efficiency, as well as rank adaptiveness. While many algorithms have been developed to optimize some of these aspects, practical ID algorithms proficient in all aspects remain absent. To fill in the gap, we introduce robust blockwise random pivoting (RBRP) that is asymptotically fast, hardware-efficient, and rank-adaptive, providing accurate skeletons and interpolation matrices comparable to the best existing ID algorithms in practice. Through extensive numerical experiments on various synthetic and natural datasets, we demonstrate the appealing empirical performance of RBRP from the aforementioned perspectives, as well as the robustness of RBRP to adversarial inputs.
@@
Randomized Numerical Linear Algebra, Interpolative Decomposition, Column Subset Selection, Adaptive Sampling
@@ 1 Introduction
The interpolative decomposition (ID) is a special type of low-rank approximation where a basis for the row/column space of a given target matrix is formed by subselecting rows/columns. This allows an ID approximation to identify structure and essential information within data. The ID has found a wide range of applications in numerical analysis, data compression, and machine learning. For example, in fast direct solvers for linear elliptic partial differential equations (PDEs), the ID is a widely used method for compressing certain off-diagonal blocks of rank-structured matrices [51, 32, 39, 40, 13, 53]. Another emerging application of the ID lies in the low-rank adaptation (LoRA) [42] for parameter-efficient fine-tuning of large language models [41, 47] where ID of weight matrices can be leveraged as good initialization for the low-rank adapter matrices [63]. More broadly, the idea of ID naturally connects to various related problems like CUR decomposition, kernel quadrature, optimal experimental design, data selection, active learning, and model pruning [49, 29, 57, 3, 15, 24, 4, 48].
Given $\mathbf{X}=\left[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\right]^{T}\in\mathbb{R}^{n\times d}$ consisting of $n$ data points $\left\{\mathbf{x}_{i}\in\mathbb{R}^{d}\right\}_{i\in[n]}$, along with a constant $\epsilon>0$ and a target rank $r\in\mathbb{N}$, we aim to construct a $(r,\epsilon)$-ID of $\mathbf{X}$:
(1) $\underset{n\times d}{\mathbf{X}}\approx\underset{n\times k}{\mathbf{W}}\ \underset{k\times d}{\mathbf{X}_{S}}\quad\emph{s.t.}\quad\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}\leqslant(1+\epsilon)\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{r}\right\|_{F}^{2}$
where $k$ is the (unknown) target rank; ${\langle\mathbf{X}\rangle}_{r}$ is the optimal rank-$r$ approximation of $\mathbf{X}$; $S=\left\{s_{1},\cdots,s_{k}\right\}\subset[n]$ are the skeleton indices corresponding to the row skeleton submatrix $\mathbf{X}_{S}=[\mathbf{x}_{s_{1}},\cdots,\mathbf{x}_{s_{k}}]^{\top}\in\mathbb{R}^{k\times d}$; and $\mathbf{W}\in\mathbb{R}^{n\times k}$ is an interpolation matrix that (nearly) minimizes the error $\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}$ for the given $\mathbf{X}_{S}$. While we focus on row-wise ID throughout this work, the column-wise ID can be defined and conducted analogously by considering the transpose of $\mathbf{X}$.
The construction of an ID can be divided into two stages:
@@
    Given $S$, compute an interpolation matrix $\mathbf{W}\in\mathbb{R}^{n\times\left|S\right|}$ efficiently111 Although (2) directly provides $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}$ as the minimizer of $\min_{\mathbf{W}}\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)=\mathcal{E}_{\mathbf{X}}\left(S\right)$, the explicit computation of $\mathbf{X}\mathbf{X}_{S}^{\dagger}$ takes $O(ndk)$ time (e.g. via (6)), which is undesired. Therefore, whenever available, information from the skeleton selection process is usually leveraged to evaluate/approximate $\mathbf{W}$ efficiently in $o(ndk)$ time (as elaborated in Section 5), which may lead to the gap $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)>\mathcal{E}_{\mathbf{X}}\left(S\right)$. that (approximately) minimizes the interpolation error: $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)\triangleq\left\|\mathbf{X}-\mathbf{W}\mathbf{X}_{S}\right\|_{F}^{2}$.
In this work, we explore fast and accurate ID algorithms from both the skeleton subset selection and interpolation matrix construction perspectives, starting with the question:
How do we construct a $(r,\epsilon)$-ID efficiently with as small a skeleton $S$ as possible,
@@ 1.1 What are Fast and Accurate ID Algorithms?
    Skeleton complexity measures the utility of skeletons selected by an ID algorithm in terms of the minimum possible rank of a $(r,\epsilon)$-ID for any given $\mathbf{X}$. Precisely, the skeleton complexity is a property of an ID algorithm defined as the minimum rank $k$ such that for any $\mathbf{X}$, $\left|S\right|=k$ is sufficient to guarantee either a $(r,\epsilon)$-ID for a deterministic algorithm or $\mathbb{E}[\mathcal{E}_{\mathbf{X}}\left(S\right)]\leqslant(1+\epsilon)\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{r}\right\|_{F}^{2}$ for a randomized algorithm.
@@
    Hardware efficiency refers to the implementation advantage that dominant cost of skeleton selection in an ID algorithm appears as matrix-matrix, instead of matrix-vector, multiplications. On modern processors, Level-3 BLAS, or matrix-matrix, operations can leverage memory cashing much more efficiently than Level-2 or Level-1 BLAS operations [7, 34]. As a result, a sequence of matrix operations implemented using Level-3 BLAS can attain significantly higher performance than an optimal implementation using Level-2 or Level-1 BLAS, especially on parallel processors.
@@
    Rank-adaptiveness refers to the ability to evaluate the skeletonization error efficiently on the fly without requiring prior knowledge of the target rank $k$.
    Remark 1.1.
    An ID algorithm is rank-adaptive if after selecting a skeleton subset $S$, it can evaluate $\mathcal{E}_{\mathbf{X}}\left(S\right)$ efficiently with at most $O\left(n\right)$ operations. That is, given any relative error tolerance $\tau>0$, a rank-adaptive ID algorithm can determine an appropriate rank $k$ with negligible (i.e. no more than $O(nk)$) additional cost, while selecting a skeleton subset $S$ of size $k$ such that $\mathcal{E}_{\mathbf{X}}\left(S\right)\leqslant\tau\left\|\mathbf{X}\right\|_{F}^{2}$.
    For the sake of analysis, we define a useful constant associated with $\mathbf{X}$:
    (3) $\eta_{r}\triangleq{\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{r}\right\|_{F}^{2}}/{\left\|\mathbf{X}\right\|_{F}^{2}},$
    which quantifies the relative tail weight of $\mathbf{X}$ with respect to the rank $r$ and represents the relative optimal rank-$r$ approximation error of $\mathbf{X}$. Notice that when $\tau<(1+\epsilon)\eta_{r}$, the algorithm outputs a $(r,\epsilon)$-ID.
@@
    ID-revealing property characterizes if the skeleton selection stage of an ID algorithm contains sufficient information for efficient construction of the interpolation matrix.
    Definition 1.1 (ID-revealing properties).
    We say a skeleton selection algorithm is
    • ID-revealing if it contains sufficient information in addition to $S$ such that the optimal interpolation matrix $\mathbf{W}=\mathbf{X}\mathbf{X}_{S}^{\dagger}$ can be evaluated efficiently in $O(nk^{2})$ time;
    • $\gamma$-ID-revealing for some $\gamma>1$ if it can construct a suboptimal interpolation matrix $\mathbf{W}\approx\mathbf{X}\mathbf{X}_{S}^{\dagger}$ with $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)\leqslant\gamma\mathcal{E}_{\mathbf{X}}\left(S\right)$ in $O(nk^{2})$ time; and
    • non-ID-revealing if neither the optimal interpolation matrix nor its approximations can be constructed in $O(nk^{2})$ time, i.e., the interpolation matrix can only be constructed by solving the least-squares problem (2) in $O(ndk)$ time.
@@ 1.2 How to Combine Adaptiveness and Randomness for ID?
Adaptiveness and randomness are two key algorithmic properties for skeleton selection. On the one hand, with adaptiveness, skeleton selection in each step is aware of the previous selections so that redundant skeleton selection can be better avoided. On the other hand, randomness balances the exploitation with exploration, thereby improving the robustness to adversarial inputs.
@@
    Greedy pivoting is a classical strategy in numerical linear algebra that involves only adaptiveness. As an example of greedy pivoting, column-pivoted QR (CPQR) [33, Section 5.4.2] starts with the input matrix $\mathbf{X}$ as the active submatrix, picks the row with the maximum norm in each step, and adaptively updates the active submatrix via projection onto the orthogonal complement of the skeletons. CPQR is rank-adaptive and ID-revealing, and usually provides low skeleton complexities in practice. However, deterministic greedy pivoting algorithms like CPQR are inherently sequential and vulnerable to adversarial inputs where the skeleton complexity can approach the problem size $n$ [43, 8].
@@
    Sampling comprises the well-researched body of skeleton selection methods that involve only randomness. Some common examples related to ID include squared-norm sampling [30], leverage score sampling [49], and DPP sampling [5, 45, 19]. As a trade-off between the skeleton complexity and efficiency, the fast ($O(nd)$-time) sampling methods like squared-norm sampling tend to suffer from high skeleton complexities that depend heavily on the matrix [20, 8]; whereas constructions of the more sophisticated distributions like leverage score and DPP are usually expensive [27, 6, 16]. Moreover, for ID, sampling methods generally fail to be rank-adaptive or ID-revealing.
@@
    Random pivoting combines adaptiveness and randomness by replacing the greedy selection of maximum norm in CPQR with random sampling according to the squared-norm distribution associated with the residual [20, 21, 8]. Closely related to the ID problem considered in this work, the idea of random pivoting has been revitalized by the inspiring recent work [8] in the context of column Nyström approximation with a nearly optimal skeleton complexity guarantee [8, Theorem 5.1]. However, analogous to CPQR, although random pivoting enjoys the desired rank-adaptiveness and ID-revealing properties, the sequential nature of random pivoting compromises its efficiency in practice.
@@
    Sketchy pivoting combines adaptiveness and randomness alternatively by first sketching $\mathbf{X}$ from right and then selecting skeletons through greedy pivoting on the sketched matrix [62, 23]. In contrast to random pivoting, the cost of sketchy pivoting is dominated by the embarrassingly parallelizable sketching process. As a trade-off for superior empirical efficiency, sketchy pivoting sacrifices the rank-adaptiveness since sketching requires prior knowledge of the target $k$ or its overestimation. Furthermore, sketchy pivoting is not ID-revealing due to the loss of information during sketching. As a simple but effective remedy for the loss of accuracy, we observe that multiplicative oversampling can remarkably improve the gap between the interpolation and skeletonization error $\mathcal{E}_{\mathbf{X}}\left(\mathbf{W}\mid S\right)-\mathcal{E}_{\mathbf{X}}\left(S\right)$, which we refer to as oversampled sketchy ID (OSID). In Section 5.2, we show that with sufficiently large oversampling, OSID is $\gamma$-ID-revealing for some small constant $\gamma>1$.
@@
Table 1: Summary of performance for skeleton selection methods that leverage adaptiveness and/or randomness. Recall from (3) that $\eta_{r}\triangleq{\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{r}\right\|_{F}^{2}}/{\left\|\mathbf{X}\right\|_{F}^{2}}$. For randomized methods, ($\mathbb{E}$) denotes that $\mathbb{E}[\mathcal{E}_{\mathbf{X}}\left(S\right)]\leqslant(1+\epsilon)\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{r}\right\|_{F}^{2}$ holds in expectation. The skeleton complexities in $(\cdot)^{*}$ are conjectured based on extensive empirical evidence without formal proofs. In the “Dominant Cost” column (showing both the asymptotic complexity and hardware efficiency), “m-v” stands for matrix-vector multiplications, which are significantly less efficient than “m-m”—matrix-matrix multiplications [34].
@@
Skeleton Selection Skeleton Complexity Dominant Cost Rank-adaptive ID-revealing Greedy pivoting/CPQR ([33, Section 5.4.2]) $k\geqslant\left(1-\left(1+\epsilon\right)\eta_{r}\right)n$ $O(ndk)$ m-v Yes Yes Squared-norm sampling ((5[30]) ($\mathbb{E}$) $k\geqslant\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ $O(nd)$ m-m No No Random pivoting (Algorithm 1 [8, 20, 21]) ($\mathbb{E}$) $k\geqslant k_{\text{RP}}=\frac{r}{\epsilon}+$ $r\min\left\{\log\left(\frac{1}{\epsilon\eta_{r}}\right),1+\log\left(\frac{2^{r}}{\epsilon}\right)\right\}$ $O(ndk)$ m-v Yes Yes Sketchy pivoting ([23, Algorithm 1][62]) ($\mathbb{E}$) $\left(k\gtrsim k_{\text{RP}}\right)^{*}$ $O(ndk)$ m-m No No RBRP (this paper) (Algorithm 2) ($\mathbb{E}$) $\left(k\gtrsim k_{\text{RP}}\right)^{*}$ $O(ndk)$ m-m Yes Yes
@@
To fill in a critical missing piece in the family of existing methods in Table 1 that leverage adaptiveness and randomness, in this work, we introduce robust blockwise random pivoting (RBRP, Algorithm 2)—an ID algorithm that is $O(ndk)$-time, hardware-efficient, rank-adaptive, and ID-revealing, with comparable skeleton complexities to the nearly optimal random pivoting [8] in practice.
@@
In particular, the plain blockwise random pivoting (e.g. an extension of the blocked RPCholesky algorithm introduced in [8, Algorithm 3]) tends to suffer from unnecessarily large skeleton complexity (cf. Figure 1 (left)) under adversarial inputs (e.g. Example 4.1) due to the lack of local adaptiveness within each block. As a remedy, RBRP leverages robust blockwise filtering (Remark 4.2)—applying CPQR to every small sampled block locally and discarding the potentially redundant points through a truncation on the relative residual of the CPQR. By choosing a reasonable block size, such robust blockwise filtering effectively resolves the inefficiency in skeleton complexity encountered by the plain blockwise random pivoting (cf. Figure 1 (right)), with negligible additional cost.